# LLMs generate structurally realistic social networks but overestimate political homophily

### v1

@@ 5 Results
Figure 2: Generated social networks from different prompting methods: Global (top), Local (middle), Sequential (bottom).
@@
Figure 5: Rates of homophily in our generated networks, per demographic variable. Ratios above 1 (marked by the grey line) indicate homophily, with higher ratios indicating more homophily.
@@
We experiment with the following LLMs: OpenAI’s GPT-3.5 Turbo and GPT-4o Brown et al. (2020); OpenAI et al. (2023), Meta’s Llama 3.1 (8B and 70B) Touvron et al. (2023), and Google’s Gemma 2 (9B and 27B) Gemma Team et al. (2024), representing a range of models across companies and sizes. We find that GPT-3.5 Turbo performs the best at matching the real social networks, so we report results from GPT-3.5 Turbo in the main text, but we report results from all models in Appendix A. We show that our main results about political homophily being most emphasized and overestimated hold for all six models. We also include sensitivity analyses, with different temperatures (Figure 13) and minor changes to the prompt (Figure 14), and show that results are stable.
We sample $N=50$ personas and we use the same set of personas for all LLM experiments. In Table 7, we report the demographic make-up of these 50 sampled personas. For each prompting method, we generate 30 networks, to capture variation in prompting (e.g., order of personas) and model response. For the graph-level metrics, we visualize their mean and standard error, along with individual data points (Figures 3 and 5). Visualizing both the standard error and individual data points capture inferential uncertainty and outcome variability, as recommended by Zhang et al. (2023).
@@ 5.1 Evaluating network structure
Thus, LLMs produce more realistic social networks when we assign the LLM to act as one persona at a time, instead of prompting it to produce the entire network at once. This is interesting, since the LLM has strictly less information under the one-persona setting: in the Local setting, it has no access to any network information, only making local decisions per persona based on demographics, and in the Sequential setting, it only knows the network based on previous personas’ choices without any ability to see into the future. In comparison, the Global method allows the LLM to take into account the entire network at once, along with all personas’ demographic information, so that it can theoretically consider dependencies between all these pieces of information. However, the LLM appears overwhelmed instead by the level of information, and produces far less realistic networks.
@@
Thus, the Sequential method is able to match real social networks on many structural characteristics. In Table 3, we quantify how well each LLM method matches the real networks, and compare to classical network models, including random graph models Erdös and Rényi (1959), small-world models Watts and Strogatz (1998), and preferential attachment models Barabási and Albert (1999). Even in the best case—when we allow the classical models to choose parameters based on the real networks—these models cannot match all of the characteristics as well as our Local or Sequential methods can. For example, when using the Kolmogorov-Smirnov statistic (Eq. 5) to measure distance between distributions, Sequential achieves an average distance of 0.330, while the best classical model (small-world) yields an average distance of 0.499, 51% higher than Sequential.
@@ 5.2 Evaluating homophily
We measure homophily for gender, race/ethnicity, religion, and political affiliation, using the ratio of observed-to-expected same-group relations (1). [^4] We cannot evaluate age homophily using Eq. 1, since “same-group” is ill-defined for a numerical variable like age, but we show in Appendix B that there is clear age homophily in the generated networks (Figure 10).
@@
In Figure 5, we show that, across all prompting methods and demographic variables, the generated networks clearly exhibit homophily. Furthermore, we see different levels of homophily for different demographic variables. For the more realistic Local and Sequential methods, homophily is by far the strongest for political affiliation: observed same-party relations are 85% more frequent than expected under Local and 68% more frequent than expected under Sequential. Rates of political homophily are even higher for the five other LLMs that we test, most extremely for GPT-4o and Llama 3.1 70B, where 100% of edges are same-party and the network fractures into two disconnected components (Appendix A).
Does this mean that LLMs actually pay the most attention to political affiliation when choosing social ties? Despite political homophily being the strongest, this could be due to correlations between political affiliation and other demographics; for example, hypothetically, if all Democrats had the same gender and race, and all Republicans had the same gender and race, then apparent homophily in political affiliation could actually be due to similarity in other demographics. To test this, we try shuffling the demographics, so that, while maintaining the same numbers of each group per demographic, each persona is randomly assigned to a group, thus removing correlations between demographics. When we run Sequential with these shuffled personas, we find that political homophily remains by far the strongest, 30 points above the next demographic variable (Figure 5), demonstrating that the LLM is, in fact, paying most attention to political affiliation when choosing social ties. We also test this hypothesis directly by prompting the LLM to generate a short reason for each friend that it selects. Similar politics is mentioned 87% of the time, while the next most-mentioned demographic variable, religion, is only mentioned 43% of the time (details in Appendix B).
LLMs overestimate political homophily.
@@
Same-group ratio (Eq. 1) 1.851 (0.016) 1.685 (0.024) Halberstam and Knight (2016) Twitter 1.404
Isolation index (Eq. 6) 0.720 (0.020) 0.530 (0.027) Halberstam and Knight (2016) Twitter 0.403
@@
Polarization (Eq. 7) 0.639 (0.037) 0.515 (0.041) Garimella and Weber (2017) Twitter, follow 0.33-0.42
@@
Table 1: Comparing political homophily in our generated networks to real-world networks. For Local and Sequential, we report the mean and standard error (in parentheses) over each method’s 30 generated networks.
@@
Given the LLM’s emphasis on political homophily, we seek to compare its level of political homophily to reported levels from real-world social networks. We are not able to compare to the eight social networks from Section 5.1 here since we do not have demographic features per node. However, we are able to find reported political homophily in several papers, covering both online and offline social networks. In Table 1, we summarize these comparisons, showing that Local and Sequential consistently overestimate political homophily across different measures of homophily.
@@
For example, Halberstam and Knight (2016) analyze political homophily on Twitter. In their data, same-party relations only appear 40.4% more often than expected, significantly lower than what the LLM predicts. They also compute the isolation index, which is the difference in average conservative exposure between conservatives and liberals (Eq. 6), with larger indices indicating greater isolation. They find an isolation index of 0.403, while the LLM’s is far higher, at 0.720 for Local and 0.530 for Sequential. Furthermore, the authors note that homophily could be overestimated in their data, since they selected users who follow politicians, which “may tend to disproportionately include individuals with strong preferences for linking to like-minded users.” Even though their homophily results could be overestimates, the LLM’s estimates still significantly exceed theirs.
@@
Isolation indices are even lower in Gentzkow and Shapiro (2011), who study ideological segregation in social networks. In face-to-face interactions, the highest isolation indices they report are 0.243 (family), 0.303 (people you trust) and 0.394 (among people who discuss politics). Finally, we also compare to Garimella and Weber (2017), who study political polarization on Twitter and define it as $p_{i}=2\cdot|0.5-\alpha/(\alpha+\beta)|$, where $\alpha$ and $\beta$ indicate how many left-leaning and right-learning users, respectively, are followed by user $i$, and they take the average $p_{i}$ over users (Eq. 7). Their measure captures the difference between observed leaning and a balanced leaning of $0.5$, with higher numbers indicating greater polarization. Using this measure, we find polarization levels of 0.639 for Local and 0.515 for Sequential, while Garimella and Weber (2017) report substantially lower numbers, with 0.33-0.42 for follows and 0.37-0.41 for retweets.
@@ 5.3 Incorporating interests
A natural question is whether demographic homophily is exaggerated because we only give the LLM demographic information, without other important details such as the person’s interests. Thus, we run an additional set of experiments where we allow the LLM to also generate interests for each persona. To generate interests, we prompt the LLM with, “In 8-12 words, describe the interests of someone with the following demographics” (full prompt in Figure 16). In Table 2, we provide examples of the generated interests, with the full list of personas with interests available online.
@@
As shown in Figure 5, by comparing Sequential vs. Sequential with interests, we find that incorporating interests increases homophily for some demographics (clearly for gender and slightly for political affiliation) and decreases homophily for the others (race/ethnicity and religion). Notably, after these adjustments, political homophily remains the strongest across the demographic variables. Thus, our finding that LLMs prioritize political homophily is robust to incorporating interests.
In fact, the interests themselves strongly encode political homophily, with evidence of political stereotyping. For example, among the Democrat personas, the most common interests are “social justice” (62.5% of Democrats vs. 0.2% of Republicans), “community service” (29.3% vs 13.9%), and “progressive policies” (18.6% vs. 0%). In contrast, among the Republican personas, the most common interests are “conservative politics” (41.6% of Republicans vs. 0% of Democrats), “church activities” (32.1% vs 13.2%), and “gardening” (23.2% vs 16.4%). We provide the complete list of top 10 interests per group in Table 5.
@@
We also analyze interests by mapping them to text embeddings, using OpenAI’s text-embedding-3-small model. In Figure 6, we visualize the embeddings, coloring them by their persona’s gender (left) and political affiliation (right). We compare these two demographic variables, since both have two, approximately equally sized groups in our sample of 50 personas. From this comparison, we can see how much more distinct the political groups are than the gender groups, demonstrating the level of political homophily encoded in the interests. Finally, we conduct a supplementary experiment where we try network generation with only interests and no demographics. As expected, homophily decreases across all demographic variables, but political homophily remains the strongest, 20 points above the next demographic variable (Figure 12).
These results demonstrate how the LLM’s emphasis on political affiliation also appears in interest generation, and, as a result, incorporating LLM-generated interests cannot help to reduce overestimated political homophily. These results build on prior results showing that LLM sometimes exhibit bias in political settings Cheng et al. (2023b); Santurkar et al. (2023); Wang et al. (2024), exploring these issues through a novel lens of homophily and social network generation.
@@ 6 Discussion
Our work has established several findings. First, “local” prompting methods result in more realistic social networks than “global” methods, even though local settings receive less information about the existing network. Second, within local methods, adding network information (i.e., Sequential) helps the LLM produce more realistic networks, so that long-tailed degree distributions are captured. Third, the LLM clearly exhibits homophily across five key demographic variables, but political homophily dominates, to the extent that it is overestimated relative to real-world measures. Finally, incorporating LLM-generated interests does not reduce these overestimates, since the interests themselves encode strong political homophily.
@@
Our findings demonstrate the promise of generating social networks with LLMs, as they are zero-shot, flexible, and able to simultaneously match many structural characteristics of real social networks. However, more needs to be done to address potential biases in social network generation with LLMs, especially with regards to political homophily. There are also other limitations to LLM social network generation: for example, while we have shown that Sequential can approximately match the means of the real networks’ metrics, its generated networks consistently demonstrate less variance than the real networks (Figure 3), reflecting known issues of LLMs to flatten demographic groups Wang et al. (2024) and lack output diversity Kirk et al. (2023). Furthermore, all of our methods require listing $N$ personas per prompt, which becomes infeasible with larger networks, due to context windows and cost. In Table 6, we conduct a big-O analysis of how the number of tokens scales with network size, revealing a tradeoff between the improved realism of Sequential and Local versus the lower costs of Global.
In future work, we hope to develop LLM network generation methods that can balance realism and costs, as well as generate networks with greater variance. We also hope to explore methods to mitigate the political bias in generated networks. Instead of LLM-generated interests, researchers could handcraft interests with fewer political stereotypes, although other methods would be needed to scale such efforts for large social networks. We also discuss in the Appendix that the LLM’s emphasis on political homophily seems to come at the cost of underestimating other types of homophily, especially for race/ethnicity, and we hope to investigate homophily along these other dimensions more systematically in future work.

### v2

@@ 5 Results
Figure 2: Examples of social networks generated by our three prompting methods: Global (top), Local (middle), and Sequential (bottom).
@@
Figure 5: Rates of homophily in our generated networks, per demographic variable. Ratios below 1 (marked by the grey line) indicate homophily, with lower ratios indicating more homophily. We visualize mean and standard error (in black) and individual data points corresponding to each network.
@@
We experiment with the following LLMs: OpenAI’s GPT-3.5 Turbo and GPT-4o (Brown et al. 2020; OpenAI et al. 2023), Meta’s Llama 3.1 (8B and 70B) (Touvron et al. 2023), and Google’s Gemma 2 (9B and 27B) (Gemma Team et al. 2024). These six models represent a range across companies, different model sizes, and proprietary (the GPT models) versus open-source (the Llama and Gemma models). We find that GPT-3.5 Turbo performs the best at matching the real social networks, so we report results from GPT-3.5 Turbo in the main text, but we report results from all models in Appendix A. We show that our main results about political homophily being most emphasized and overestimated hold for all six models. We also include sensitivity analyses, with different temperatures (Figure B5) and minor changes to the prompt (Figure B6), and show that results are stable.
We sample $N=50$ personas and we use the same set of personas for all LLM experiments. In Table C1, we report the demographic make-up of these 50 sampled personas. For each prompting method, we generate 30 networks, to capture variation in prompting (e.g., order of personas) and model response. For the graph-level metrics (all metrics in Section 4 besides degree distribution), we visualize their mean and standard error, along with individual data points (Figures 3 and 5). Visualizing both the standard error and individual data points capture inferential uncertainty and outcome variability, as recommended by Zhang et al. (2023). For degree (Figure 4), we visualize the pooled degree distribution over nodes in the networks, as described in Section 4.
@@ 5.1 Evaluating network structure
Thus, LLMs produce more realistic social networks when we assign the LLM to act as one persona at a time, instead of prompting it to produce the entire network at once. This is interesting, since the LLM has strictly less information under the one-persona setting: in the Local setting, it has no access to any network information, only making local decisions per persona based on demographics, and in the Sequential setting, it only knows the network based on previous personas’ choices without any ability to see into the future. In comparison, the Global method allows the LLM to take into account the entire network at once, along with all personas’ demographic information, so that it can theoretically consider dependencies between all these pieces of information. However, the LLM is not able to effectively leverage all of these dependencies and instead produces far less realistic networks.
@@
Thus, the Sequential method is able to match real social networks on many structural characteristics. In Table B1, we quantify how well each LLM method matches the real networks, and compare to classical network models with few parameters, including random graph models (Erdős and Rényi 1959), small-world models (Watts and Strogatz 1998), and preferential attachment models (Barabási and Albert 1999). Even in the best case—when we allow the classical models to choose parameters based on the real networks—these models cannot match all of the characteristics as well as our Local or Sequential methods can. For example, when using the Kolmogorov-Smirnov statistic (Eq. 7) to measure the distance between the generated versus real networks’ distributions per structural characteristic, Sequential achieves an average distance of 0.330, while the best classical model (small-world) yields an average distance of 0.499, 51% higher than Sequential. Thus, the LLM achieves better realism than these classical network models when generating social networks, while maintaining their advantages of requiring no additional training or minimal need to define parameters.
@@ 5.2 Evaluating homophily
We measure homophily for gender, age, race/ethnicity, religion, and political affiliation, using the ratio of observed-to-expected proportion of cross-group edges (Eq. 1).
@@
In Figure 5, we show that, across all prompting methods and demographic variables, the ratio is significantly below 1, indicating that the generated networks clearly exhibit homophily. Furthermore, we see different levels of homophily for different demographic variables. For the more realistic Local and Sequential methods, homophily is by far the strongest for political affiliation: observed cross-party relations are 82% less frequent than expected under Local and 66% less frequent than expected under Sequential. Rates of political homophily are even stronger for the other LLMs that we test, most extremely for GPT-4o and Llama 3.1 70B, where none of the edges are cross-party and the network fractures into two disconnected components (Appendix A).
Does this mean that LLMs actually pay the most attention to political affiliation when choosing social ties? Despite political homophily being the strongest, this could be due to correlations between political affiliation and other demographics; for example, hypothetically, if all Democrats had the same gender and race, and all Republicans had the same gender and race, then apparent homophily in political affiliation could actually be due to similarity in other demographics. To test this, we try shuffling the demographics, so that, while maintaining the same numbers of each group per demographic, each persona is randomly assigned to a group, thus removing correlations between demographics. When we run Sequential with these shuffled personas, we find that political homophily remains by far the strongest (Figure 5), demonstrating that the LLM is, in fact, paying most attention to political affiliation when choosing social ties. As an additional test, we also try ablations of the demographics, where we present the LLM with only one demographic variable at a time, or with two variables, one being political affiliation and the other being one of the four others. Here, we also find that political affiliation continues to be the dominant factor: when only one variable is presented, levels of homophily increase for all demographics but it remains the highest for political affiliation; when two variables are presented, political homophily is always stronger than the other demographic’s homophily (Figure B2).
Demographic Reason %
Political affiliation 86.7%
Religion 43.0%
Age 21.8%
Race/ethnicity 12.1%
Gender 7.3%
Table 1: Frequency that each demographic is part of the LLM’s reason for choosing a persona as a friend.
@@
Cross-group ratio (Eq. 1) $\downarrow$ 0.180 (0.015) 0.340 (0.022) Halberstam and Knight (2016) Twitter 0.528
Same-group ratio (Eq. 8) $\uparrow$ 1.851 (0.016) 1.685 (0.023) Halberstam and Knight (2016) Twitter 1.404
Isolation index (Eq. 9) $\uparrow$ 0.729 (0.020) 0.530 (0.027) Halberstam and Knight (2016) Twitter 0.403
@@
Polarization (Eq. 10) $\uparrow$ 0.639 (0.037) 0.515 (0.041) Garimella and Weber (2017) Twitter, follow 0.33-0.42
@@
Table 2: Comparing political homophily in our generated networks to real-world networks. We consider different measures of homophily and indicate with $\downarrow$ or $\uparrow$ the direction that indicates greater homophily for that measure. For Local and Sequential, we report the mean and standard error (in parentheses) over each method’s 30 generated networks.
Finally, we directly test what the LLM is paying attention to by prompting it to generate a short reason for each friend that it selects. Then, given the reason, we use GPT-4o to classify the reason, e.g., “I’m a woman too and we share the same religion and political affiliation” is classified as [gender, religion, political affiliation]. We allow the LLM to generate free-text reasons during network generation since we do not want to constrain its response (e.g., if it is using other information, such as degree or ID). We find that political affiliation strongly dominates here as well: it is part of the reason 86.7% of the time, while the next most-mentioned demographic, religion, is only mentioned 43.0% of the time (Table 1). As a caveat, prior work has shown that we cannot always trust an LLM’s own explanation for its choices (Agarwal, Tanneru, and Lakkaraju 2024). However, given the alignment of these results with our other results, there is strong evidence that LLMs pay the most attention to political affiliation when generating social relations.
LLMs overestimate political homophily.
@@
Given the LLM’s emphasis on political homophily, we seek to compare its level of political homophily to reported levels from real-world social networks. We are not able to compare to the eight social networks from Section 5.1 here since we do not have demographic features per node. However, we are able to find reported political homophily in several papers, covering both online and offline social networks. In Table 2, we summarize these comparisons, showing that Local and Sequential consistently overestimate political homophily across different measures of homophily. For example, Halberstam and Knight (2016) analyze political homophily on Twitter. In their data, cross-party relations appear 47% less often than expected, which indicates homophily, but not as strong as what the LLM predicts. In addition to the cross-group ratio, we can also compute a same-group ratio, using the ratio of observed-to-expected proportion of same-group edges, where a higher ratio indicates more homophily (Eq. 8). Using this measure, we find that same-party relations only appear 40% more often than expected in the Twitter data, while same-party relations are 85% and 68% more frequent than expected for Local and Sequential, respectively.
@@
Halberstam and Knight (2016) also compute the isolation index, which is the difference in average conservative exposure between conservatives and liberals (Eq. 9), with larger indices indicating greater isolation. They find an isolation index of 0.403, while the LLM’s is far higher, at 0.720 for Local and 0.530 for Sequential. Furthermore, the authors note that homophily could be overestimated in their data, since they selected users who follow politicians, which “may tend to disproportionately include individuals with strong preferences for linking to like-minded users.” Even though their homophily results could be overestimates, the LLM’s estimates still significantly exceed theirs.
@@
Isolation indices are even lower in Gentzkow and Shapiro (2011), who study ideological segregation in social networks. In face-to-face interactions, the highest isolation indices they report are 0.243 (family), 0.303 (people you trust) and 0.394 (among people who discuss politics). Finally, we also compare to Garimella and Weber (2017), who study political polarization on Twitter and define it as $p_{i}=2\cdot|0.5-\alpha/(\alpha+\beta)|$, where $\alpha$ and $\beta$ indicate how many left-leaning and right-learning users, respectively, are followed by user $i$, and they take the average $p_{i}$ over users (Eq. 10). Their measure captures the difference between observed leaning and a balanced leaning of $0.5$, with higher numbers indicating greater polarization. They report polarization levels of 0.33-0.42 for following relations and 0.37-0.41 for retweets (increasing over time), while we find significantly higher levels in the LLM’s generated networks, with 0.639 for Local and 0.515 for Sequential.
@@
Figure 6: Ratio of observed-to-expected proportion of edges, for all pairs of demographic groups (Eq. 2). All subfigures share the same colormap (right). Groups with at least 10 nodes are kept.
Homophily between pairs of groups.
As described in Eq. 1, we measure homophily as the ratio of observed-to-expected proportion of cross-group edges. Now, we extend this definition to compute an observed-to-expected ratio for any pair of groups, $A$ and $B$ (where $A$ could equal $B$):
$H_{AB}$ $=\frac{\frac{\sum_{ij}A_{ij}\cdot\mathds{1}[g_{i}=A]\cdot\mathds{1}[g_{j}=B]}{E}}{\frac{N_{A}(N_{B}-\mathds{1}[A=B])}{N(N-1)}}.$ (2)
This measures the observed proportion of edges that are between nodes in groups $A$ and $B$ divided by the expected proportion of edges between nodes in these groups. In Figure 6, we visualize $H_{AB}$ for all pairs, for each of the five demographic variables. For each variable, we keep all groups with at least 10 personas in our set of 50 personas (Table C1), and compute $H_{AB}$ for each pair of groups, reporting the mean ratio over the 30 generated networks from Sequential.
We find, as expected, that the diagonal (i.e., same-group ratios) tends to be above 1, although notably it is not for relations between men, with a ratio of 0.99, while relations between women have a ratio of 1.43. Thus, even within one demographic variable, the LLM’s levels of homophily vary for different groups, such as homophily within women being stronger than homophily within men. Variability across groups is present for all demographic variables: most extremely, for religion, relations between Catholics occur 89% more frequently than expected, while relations are only 5% more frequent for personas that identify as Unreligious. This plot also reveals that not all cross-group relations are equally unlikely: for example, when we divide age into three age groups, the adjacent age groups have higher cross-group ratios (0.98 for under 30 and 30-59 and 0.97 for 30-59 and 60+) compared to the non-adjacent age groups (0.87 for under 30 and 60+). Finally, this plot reiterates that political affiliation has the lowest cross-group ratio, at 0.34, and reveals that it also has among the highest same-group ratios, although the LLM gives Democrats stronger same-group preferences, with a same-group ratio of 1.85, than Republicans, with a same-group ratio of 1.54.
@@ 5.3 Incorporating interests
A natural question is whether demographic homophily is exaggerated because we only give the LLM demographic information, without other important details such as the person’s interests. Thus, we run an additional set of experiments where we allow the LLM to also generate interests for each persona. To generate interests, we prompt the LLM with, “In 8-12 words, describe the interests of someone with the following demographics” (full prompt in Figure C2). In Table 3, we provide examples of the generated interests, with the full list of personas with interests available online.
@@
As shown in Figure 5, by comparing Sequential vs. Sequential with interests, we find that incorporating interests strengthens homophily for some demographics (clearly for gender and age and slightly for political affiliation) and weakens homophily for the others (race/ethnicity and religion). Notably, after these adjustments, political homophily remains the strongest across the demographic variables. Thus, our finding that LLMs prioritize political homophily is robust to incorporating interests.
In fact, the interests themselves strongly encode political homophily, with evidence of political stereotyping. For example, among the Democrat personas, the most common interests are “social justice” (62.5% of Democrats vs. 0.2% of Republicans), “community service” (29.3% vs. 13.9%), and “progressive policies” (18.6% vs. 0%). In contrast, among the Republican personas, the most common interests are “conservative politics” (41.6% of Republicans vs. 0% of Democrats), “church activities” (32.1% vs. 13.2%), and “gardening” (23.2% vs 16.4%). We provide the complete list of top 10 interests per demographic group in Table B4.
@@
We also analyze interests by mapping them to text embeddings, using OpenAI’s text-embedding-3-small model. In Figure 7, we visualize the embeddings, coloring them by their persona’s gender (left) and political affiliation (right). We compare these two demographic variables, since both have two, approximately equally sized groups in our sample of 50 personas. From this comparison, we can see how much more distinct the political groups are than the gender groups, demonstrating the level of political homophily encoded in the interests. Finally, we conduct a supplementary experiment where we try network generation with only interests and no demographics. Homophily decreases across all demographic variables, but political homophily remains by far the strongest (Figure B4).
These results demonstrate how the LLM’s emphasis on political affiliation also appears in interest generation, and, as a result, incorporating LLM-generated interests cannot help to reduce overestimated political homophily. These results build on prior results showing that LLM sometimes exhibit bias in political settings (Cheng, Piccardi, and Yang 2023; Santurkar et al. 2023; Wang, Morgenstern, and Dickerson 2024), exploring these issues through a novel lens of homophily and social network generation.
@@ 6 Discussion
Our work has established several findings. First, with the right prompting method, the LLM is able to simultaneously match many structural characteristics of real social networks, outperforming classical network models with few parameters. Second, “local” prompting methods produce more realistic networks than “global” methods and, within local methods, adding network information (i.e., Sequential) helps the LLM capture long-tailed degree distributions. Third, the LLM exhibits clear homophily across five key demographic variables, but political homophily dominates, to the extent that it is overestimated relative to real-world measures. Finally, incorporating LLM-generated interests does not reduce these overestimates, since the interests themselves encode strong political homophily.
Future directions.
Our findings demonstrate the promise of generating social networks with LLMs, as they are zero-shot, flexible, and structurally realistic. However, more needs to be done to address potential biases, especially with regards to political homophily. One possibility could be to incorporate more information per persona: we found in our ablations that the LLM tends to produce stronger homophily per variable when fewer variables are provided (Figure B2), so we might see political homophily reduce if more variables or richer descriptions (such as interests) are provided per personas. However, given our findings on LLM-generated interests, researchers may want to handcraft interests with fewer political stereotypes, although other methods would be needed to scale such efforts to large social networks. Furthermore, while we focus in this work on how LLMs overestimate political homophily compared to real-world measures, in the Appendix we discuss real-world measures of homophily for the other demographic variables and find that the LLM, in fact, seems to underestimate homophily for race/ethnicity and religion (Table B2), calling for future work that thoroughly investigates potential biases along the other demographic dimensions.
@@
There are also other limitations to LLM social network generation: for example, while we have shown that Sequential can approximately match the means of the real networks’ metrics, its generated networks consistently demonstrate less variance than the real networks (Figure 3), reflecting known issues of LLMs to flatten demographic groups (Wang, Morgenstern, and Dickerson 2024) and lack output diversity (Kirk et al. 2023). Furthermore, all of our methods require listing $N$ personas per prompt, which becomes infeasible with larger networks, due to context windows and cost. In Table C2, we conduct a big-O analysis of how the number of tokens scales with network size, revealing a tradeoff between the improved realism of Sequential and Local versus the lower costs of Global. To address this tradeoff, future work could explore how to make the Sequential or Local methods more scalable: for example, instead of listing $N$ personas per prompt, where each persona chooses from the full list of all other personas, one could provide a subset of the other personas. In Appendix C.2, we discuss a simple implementation of this idea where the subsets are sampled uniformly at random, and show how this extension enables us to generate networks with 300 personas (sampling 30 personas per prompt). These larger networks are similar to our smaller networks in terms of homophily and some structural characteristics, although density and clustering drop due to the subsampling (Figure C6).
In future work, we hope to further extend these methods to make them both scalable and realistic, as well as generate networks with greater variance. We also hope to study whether our results on political homophily generalize outside of the US, and to explore methods to mitigate the political biases we observe in this work.